OpenStudy (stormoshawty):

Solve y= 1/2x - 6 2x +6y=19 using substitution. @shamil98

OpenStudy (shamil98):

so substitute 2x + 6(1/2x-6) = 19

OpenStudy (shamil98):

let me know what you get for the value of x.

OpenStudy (stormoshawty):

X = 5

OpenStudy (shamil98):

2x + 3x - 36 = 19 5x = 55 x = ?

OpenStudy (stormoshawty):

x=11

OpenStudy (shamil98):

yep. now sub back in to find y 2(11) + 6y = 19

OpenStudy (stormoshawty):

\[y = -\frac{ 1 }{ 2 }\]

OpenStudy (shamil98):

22 + 6y = 19 6y = -3 y = -1/2 yep good job

OpenStudy (stormoshawty):

Thank you Shammy.

OpenStudy (shamil98):

mhm